myininaya (myininaya):

Fun Calculus Question! find numbers a and b lim sqrt{ax+b}-2 =1 x->0 ---------- x

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sqrt{ax+b}-2}{x}=1\]

OpenStudy (mathteacher1729):

How far did you get on this prob before you got stuck? :)

myininaya (myininaya):

i just started doing it. i figure it would be a fun problem for everyone to attemp since it is in the problems plus section of my calculus book

OpenStudy (mathteacher1729):

Cool. Well, I got the answer in about 2 seconds using Geogebra and some sliders. To do it analytically might take a little longer. :)

OpenStudy (anonymous):

i think i saw this somewhere before,,,,but i'm sure i never solved it...it's fun to try now :D

OpenStudy (mathteacher1729):

myininaya, if you wanna see the geogebra answer I can share a screen with you on my virtual whiteboard.

myininaya (myininaya):

sure! lets do it

myininaya (myininaya):

what did you get for a and b?

myininaya (myininaya):

a=4 b=4? is that right let me check

OpenStudy (anonymous):

im gonna join :D if you dont mind

OpenStudy (mathteacher1729):

Hope that was enjoyable, play with geogebra, it's awesome! :D

myininaya (myininaya):

ok since the bottom goes to 0 the top would have to go to 0 \[\sqrt{ax+b}-2=0\] => x->0, b=4 \[\lim_{x \rightarrow 0}a \frac{1}{2}(ax+b)^\frac{-1}{2}=\lim_{x \rightarrow 0}\frac{a}{2\sqrt{ax+b}}=\frac{a}{2\sqrt{b}}\] but b=4 \[\frac{a}{2*2}=\frac{a}{4}=1\] => a=4

OpenStudy (radar):

I like that approach myininaya

myininaya (myininaya):

:)

OpenStudy (anonymous):

somewhere kinda unseen to me, but i'm still curious how you get the second limit :)

myininaya (myininaya):

you mean a?

OpenStudy (anonymous):

a/2squr(ax+b) but how you approach to that

OpenStudy (anonymous):

L' hospital rule?

myininaya (myininaya):

yes

myininaya (myininaya):

we know the limit is suppose to be 1 so thats why we set a/4=1

OpenStudy (anonymous):

oh ok, i see it now...Great and fun :)